OpenStudy (raveninneverland):

Dealing with parallel and perpendicular lines, and their slopes. If two lines are parallel, and both of their slopes are 1.5, then I need to know what the two slopes relationship is. I first thought it was that they are both the same. On another question, I have two perpendicular lines. One line has a slope of -2/3, and the line it crosses over has a slope of 1.5. I need to know the relationship between the two slopes of these lines. Please don't just give me an answer, I would appreciate it if you told me how to find the relationship between two slopes.

OpenStudy (phi):

**I first thought it was that they are both the same.*** that sounds right. parallel lines have identical slopes for perpendicular slopes, you have \[ m_1= -\frac{2}{3} , m_2 = 1.5\] if we write 1.5 as 3/2: \[ m_1= -\frac{2}{3} , m_2 = \frac{3}{2}\] do you see any "pattern" ?

OpenStudy (raveninneverland):

Hm. Is it because they are both the same number, but one is negative? (on the perpendicular one)

OpenStudy (phi):

2/3 and 3/2 are not the same number. But they are related

OpenStudy (raveninneverland):

OH. Oh, I can see it now. (Haha, I was seeing them as the same number) Are they related because they involve the same numbers?

OpenStudy (phi):

one is the "flip" of the other. actually the negative "flip" Of course, people don't say flip, they say "negative reciprocal" if one slope is \(m_1\), the slope of its perpendicular is \( m_2= - \frac{1}{m_1} \)

OpenStudy (phi):

example \[ m_1= -\frac{2}{3} \\ m_2= - \frac{1}{-\frac{2}{3}} \\ m_2= - \frac{1\cdot -\frac{3}{2}}{-\frac{2}{3}\cdot -\frac{3}{2}}= -\frac{ -\frac{3}{2}}{1}= \frac{3}{2}\\ \]

OpenStudy (phi):

(But I think it's easier to just "flip" -2/3 to -3/2 then negate it, to get 3/2)

OpenStudy (raveninneverland):

Oh! I can understand it now. :D Thank you very much! <3